OpenStudy (babyslapmafro):

Please help. I am asked to use a total differential to approximate the change in the values of f from P to Q and then compare the estimate with the actual change in f. I know how to obtain an estimate using the total differential but how do you find the actual change in f? f(x,y)=x^2+2xy-4x P(1,2) Q(1.01,2.04)

OpenStudy (valpey):

The actual change in f is f(Q)-f(P). \[f(P) = 1^2 + 2*1*2-4*1\]\[f(Q) = 1.01^2 + 2*1.01*2.04-4*1.01\]